OpenStudy (anonymous):

A motorboat takes 5 hours to travel 450 miles going upstream. The return trip takes 3 hours going downstream. What is the rate of the boat in stil water and what is the rate of the current? Please help I've been stuck on this for soo long!

OpenStudy (anonymous):

Let x represent the rate of the boat. Let c represent the rate of the current. D = r * t Rate against current is (x-c) miles per hour. Rate with current is (x+c) miles per hour. Equation 1: 450 = (x-c)*5 Equation 2: 450 = (x+c)*3 Then solve... I would divide both sides of the first equation by 5 and both sides of the second equation by 3 and solve using the elimination method. http://www.wolframalpha.com/input/?i=solve {450%3D%28x-c%29*5%2C450%3D%28x%2Bc%29*3} Hope this helps.

OpenStudy (anonymous):

Yeah that's the way my techer taight me how to solve them and I ended up getting it wrong still, somewhre I'm making a mistake. Could I ask what you got for your answer please?

OpenStudy (anonymous):

I really need the correct answer for this to pass a final and can't get it.

OpenStudy (anonymous):

Follow the link... after dividing: 90 = x - c 150 = x+ c Add the equations together 240 = 2x divide by 2 120 = x boat 120 miles per hour... (fast for a boat?) Current 30 miles per hour...

OpenStudy (anonymous):

A 120 mph boat? WOW. But thanks for helping you're a lifesaver!

OpenStudy (anonymous):

Maybe it should be an airplane or something... 450 miles is far to go in 5 hours.

OpenStudy (anonymous):

Agreed, but that's how they had it worded. A motorboat?